OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the area of a rectangle A = LW? A/WL = 1 A/W = L A = L/W A/L = W

OpenStudy (anonymous):

A = L/W is not equivalent to A = LW it makes sense, L/W CANNOT EQUAL LW.

OpenStudy (anonymous):

the other options are just rearrangements of the formula. example: A = LW , then solve for L, so divide each side by W --> L = A/W...